DFR Telecoms Diary
5 January : Posted by Sam
Welcome to 2020
Sam and Charles met at Norchard to investigate the token circuit. Starting In the signal box, they used a meter and a BT301C to look up the line. This identified a high resistance short-circuit at approximately 200 m from the signal box. After a quick walk to some cabinets this was determined to not be a contributing factor in the failure. They instead put a loop on the line and went to Parkend to establish whether they could see the loop or not. The loop was not visible from Parkend. They went to Whitecroft and determined that the fault was north of the crossing, using the 301C tester they determine that it was approximately 800 m from the crossing. Moving onto DP1G they then established that it was at the flailed joint, a quick inspection of this revealed that it was rusty and seriously corroded, and open-circuit.
They then returned to Norchard, where they found Roger H. They looked at what parts we had to joint the cable and went to find a short length of cable to splice across the fault. A short length was repaired and is ready for transporting to site on Wednesday. Two drainpipe joints are proposed to cover the gap.
Roger and Sam then went to look at the Norchard to asterisk Junctions, one of which was failing. They had some issues as the relay set would not busy out when they moved the connector in the front. Instead they jacked the set out. This produced engaged tone when selected, which seems a better compromise than it simply failing.
They then went back to the office and had a discussion about how to run the clocks over a digital network. They decided to try and experiment using broadcast messages across the network to run a series of clocks from a master controller.
Finally They called in at Parkend to look at the sip phone in the signal box which was misbehaving. This was cured by a reset and they took the opportunity to fully swap the network switches over.
8 January : Posted by Martin
The Group held a meeting to discuss the programme of work for the coming year. Despite having retired, John still had one outstanding job to complete at Parkend Signal Box. He changed a faulty relay coil on relay 6K on the concentrator.
Notes of meeting of DFR Telephone Group in Norchard Seven and Wye room at 10:00 on 8th January 2020
Present: - Sam Bolingbroke, Charles Bristow, Allan Cantwell, Rick Gillingham, Martin Lunn, Peter Medcalf, Ray Willey.
1. First priority was determined to be re-establishing the cable connection between Norchard and Parkend by replacing damaged cable.
a. Complete the cable replacement between Whitecroft and Parkend.
i. New fifty pair cable. With two pipe joint connections. Considered not to be a good idea, joints only present problems in the future.
iii. Perhaps a new design of accessible pipe joint sunk into a concrete base is the best way to go here.
c. Reluctantly it was agreed to a two pipe joint, which was said could be done in 1 day.
d. It is important that all pairs should be tested and a record of wire colours recorded before joint sealed in pipe.
e. After an unsuccessful practice effort to strip the sheath off of a length of cable, the difficulty was caused of the low ambient temperature. It was said that a tent and heating would be made available.
i. Due to this difficulty it was thought jointing could take as long at 2 – 3 working days.
ii. After the meeting it was suggested that temporary 10 pair connection be made and the 50 pair connected when warmer weather prevailed.
iii. It was revealed that the first cable joints to be made are at DP 1E and DP1EA where new cabinets have been provided.
2. Work on the Digital system throughout the railway to be halted in favour of establishing a reliable, tried and tested system at Parkend. Digitalisation of the railways major stations and associated operational areas would take place when the system has been proved. It is possible this may take as long as three years to complete.
3. Having successfully re-established telephone cabling throughout the railway the Strowger exchanges and associated equipment can be expected to have an operational life of between 5 and 10 years or more. If and when the digital system is in a reliable working order, the service currently provided by the Norchard exchange will be gradually swapped to the Digital service.
4. Sam agreed to provide an operations plan for the rework of the digital system and keep the telecoms group up to date with developments and progress. Initial program to be provided in about two weeks, i.e. before 23rd January 2020.
5. Most of our cable problems are caused by contractors. Any costs incurred repairing such damage should be recovered from the contractor insurance. Sam is to ensure that this is the case in the future.
6. It was suggested that some form of protection from flails, tractor or diggers should be provided around DP’s and any sort of cable joint.
7. It I very important that maintenance of the current cable is maintained. This should be a priority over any digital work for some years to come.
22 January : Posted by Peter
Peter started the day by collecting the Netgear switch from Parkend. He then met Alan, Charles, Martin and Rick at the cycle cafe. After a coffee Alan, Charles and Peter departed for the junction, a bit of head scratching as to where they had left off. The data cables have all now been installed but need to be tested, The P.A. wiring needs to be finished. Peter took the transformer home to find suitable fixings, it will be mounted in the data cabinet making for a maintenance dream, easy access! Bit of a closet rant, for the people who do not think about maintaining equipment after it is installed in the wrong place. The test sockets at the junction all still need to be fitted so about a month’s work for Alan and Peter and then hopefully back to doing the wipe ups on the bench in readiness for re cabling the office. Gents please let me know your thoughts on this and for anyone who wishes a bit of hands on data cabling experience let me know. Whilst at the Junction Peter was notified, he had made a drop off, the bolts for fitting the earth cables to the box conn 501 in the signal box were too short, sorry about that one Charles. The bolts have only enough length to fix 1 eye there are 3 eyes to be fitted in the box conn. Charles fitted a new earthing strip to the exchange, a nice job.
Alan departed at lunch time. Back in the office at lunch time Martin raised a question about last weeks overview of the data cabling, another impromptu training session then ensued. If you chaps want some homework try browsing to your router, might be handy to know who has what. Charles, Peter can see the need for some sort of training budget, could we lever it out of the railway? Rick and Peter went to the exchange to find a suitable location to fit a data switch that has POE. It was agreed the switch can be nicely fitted below the gateway making a neat and tidy job. Hopefully this will be fitted Saturday.
5 February : posted by Charles
This morning Ray and Martin visited Parkend to examine the Group Selector 1 Switch that Ray had busied out last week. It proved impossible to remove the switch from the rack, so Ray and Martin prodded and poked it and eventually it started to work again and dial tone was transmitted to the handset. Allan and Charles looked at the Wickham Trolley for the first time just to get a feel for a possible restoration job to help the Telecom Group move up and down the line. There are no seats but a petrol engine, fuel tank and radiator are in place. Lynne will put us in touch with the working members who are known to have had an interest in the past.
Allan and Charles also considered how to address the Whitecroft/Parkend cable repairs by planning a considered approach to tools needed, cable excavation and re-dressing of the severed cable and insertion of a 5 yard (approx.) link.
12 February : posted by Rick
A meeting was held at Kaplans today with the telecoms group. Those in attendance - Rick Gillingham - Ray Willey - Martin Lunn - Peter Medcalf - Charles Bristow - Alan Cantrell
As Sam has not been contactable for 2 weeks and his last e mail said he did not want to manage the group any longer it was decided that we would have to manage ourselves. It was agreed that for the time being Rick would deal with Strowger problems, Peter and Roger would deal with Digital problems, Charles would deal with finance and the board and Martin would work on computer records. If anyone has any particular problems the group will discuss and decide what action should be taken.
Although the group is now dealing with Strowger and Digital it also agreed that we should not have demarcation within the group and also work towards a closer working relationship with S and T.
The really urgent problem that has to be dealt with asap is the lack of communication between Norchard, Whitecroft and Parkend caused by the catastrophic cable damage between Whitecroft and Parkend to which no real solution has been found in the last 6 months.
The meeting finished about 11 am.
Charles and Ray went home , Rick Martin Peter and Alan returned to Norchard . Rick and Peter spoke to Roger Phelps about new cable and he said he would look into it but no promises. Rick and Peter went to the office and spoke to Duncan to tell him what had been decided. He seemed to be aware that there was a problem with our Department and seemed happy that we had resolved it. There is a board meeting next Monday 17th February in the Severn and Wye room from 7 til 9.
After lunch we all went to look at the cable between Whitecroft and Parkend. Unfortunately there is no good news - the original damaged wires have been cut back so we will have to tone them out again - the cable from there to Parkend has also had the wires cut or disconnected in at least 3 separate places - there appears to be no good reason for someone to do this. We then went for a cup of tea to discuss the best way forward.
19 February : posted by Charles
This morning Martin, Allan and Charles met to discuss and plan the resources required to reinstate various communication services north of Whitecroft. In a wide-ranging attempt to identify best and minimal cost solutions, they took into account the not only the cabling, and the logistics of transporting it to site, but also trunking, excavation and cable jointing. The outcome will be shared with the Telecom group before returning to the DFR Operational Board. In the afternoon, Martin and Charles visited Parkend Box to identify and measure the access below the Axle Counter Panel for the installation of additional Telecom group equipment.
26 February : posted by Charles / Rick
Today the Carriage Group arranged a "Toolbox" learning session at lunch time about Health and Safety issues relating to hand tools. Charles was invited to attend what proved to be an excellent session well-paced and presented in a friendly fashion. The selection of the right tool for the job was emphasised, and then the use, maintenance, storage and disposal were discussed with hands-on examples of tools such as hammers, screwdrivers, files, chisels etc. It was a good example of a mutual improvement class, which deserves to be repeated in greater depth and breadth for the benefit of all workers in the different railway groups. Thank you Ollie!
Tea coffee and cake in our room at Norchard with Ray Martin Charles Allan and Rick. Then we went to Parkend - Ray returned to working on a recurring fault - the rest of us visited the new Signal box where we found Roger Phelps and had a meaningful discussion on what needs to be done to enable us to overcome our present problems and our future development. We then walked a short length of line to verify existing cabling. All of us returned to Norchard for lunch.
Ray then returned to Parkend to check everything was now ok ( all tickety boo ). Allan and Martin then searched Norchard site for an alleged reel of 50 pr cable without success and then moved on to Lydney Junction with the same result. Was this a wild goose chase? Rick spoke to Duncan about a pro forma to apply for urgently needed cable. Was advised not required and apply using Word !!
Also heard about developement at Norchard requiring moving our caboose - have arranged to meet Chris Bull for full info.
4 March : posted by Rick
Just Rick Martin Charles and Allan in today. Tea and cakes as usual and a discussion about the telecom dept. Allan and Charles went off to look at a water heater, a freezer and one of the portacabins
that had electric problems and then started the routine check of Norchard telephones.
Martin and Rick went to look at the SPT at the stop board at Norchard which was dead. When it was opened up it worked ok and when it was put back together it worked ok.
They then went to the Signal box to check and found the battery had a low volts alarm. SPT 2Z was found to be permanently calling and was disconnected. The batteries were soon back to normal. On testing SPT 2Z a battery contact was proved out to line. As the weather had deteriorated it was decided to leave finding it until another day.
Allan and Martin went home after lunch. Charles continued with the routine checks and Rick went to Norchard exchange to clear some outstanding faults. Both left at about 15.00 hrs.
7 March : posted by John
I went to Parkend to look at the multi-level relay set being used to connect to the asterisk. I gave it as good a check as I was able with my multi meter. The pulse ratio seemed fine. However I could not disconnect the set from the asterisk as the records have not been completed. I was therefore unable to connect the set to the AT5422 to measure the pulse speed and number of digits. I did not want to have to start unsoldering wires on the back of the shelf to disconnect the asterisk. The junction should go via the S&T connectors to allow a quick disconnection for testing. The regen appeared to work well and I set the output pulse speed to be a little higher in the hope that the Grandstream ATA would be happier. The regen looked dry but there was no oil for me to use to relubricate the item. I may need to return when the digital boys are on site. I was thinking of wiring some spare level final selector outlets to the S&T connectors so that the AT5422 could "dial up" the junctions for easy testing. I may need to discuss this with Rick.
11 March : posted by Rick
Those in today Rick, Martin, Charles and Allan. Started with tea/coffee and cakes (supplied by Allan's wife) and we were joined by Steve Harris ( S & T ) for a general chat and what future work is required.
Consequently it was a late start as we all went to Parkend. We removed the old cable wiring from one of the newly fitted cabinets and then tried to put the existing cable in to the cabinet but because of the cold, the cable was difficult to bend and we had to dig back until the cable could be bent sufficiently to be fitted in to the cabinet. We then visited the second cabinet and had a similar experience but did manage to fit the cable in eventually. Now we just need the new cable to be supplied and laid out along the track.
After a late lunch we returned to Parkend to look for a backboard for one of the cabinets which was in the signal box. Allan took this home to ready for fitting next week. Martin then went home while Rick and Charles returned to Norchard to have a quick look at SPT 2z as the DP records do not seem to be accurate. It looks as if it routes through DP 3R. Further time needed. A good day in some respects.
LOCK DOWN - March to July
Corona virus locked down the railway and work virtually stopped. Rick went in at intervals to do essential routine maintenance.
4 July : posted by Rick
Charles checked out the water heater in Lydney Junction Signal box. Disconnected it and took it home for repair. Peter and Rick connected the new PA transformer at the Junction and tested it working ok.
8 July : posted by Rick
Only Martin Charles and Rick in today. We supplied John Metherall with a reel of six wire and a switch key for use in the museum to allow him to set up a stop/go system to control visitors entering. After a little chat amongst ourselves and updating the latest DFR news Charles went off to look at water heaters and other electrical stuff and to liaise with Duncan about the repairs required.
Martin and Rick went to routine check Norchard SPTs. Charles later came to assist. After checking six of them - three were working ok and three were faulty. One may need a new phone and the other two were due to cable damage by other DFR staff. Wasp nests were found in two cabinets and Rick suffered two wasp stings. The SPT at Norchard crossing hut was repaired by refixing the cable to the hut and making a secure repair in a block terminal. The SPT at signal post 2 had the drop wire damaged by grass cutting and will have to wait for repair another day and will require some sort of wasp control.
During the day we met Chris Bull and vented our feelings on damage to our cables. We also discussed the problem of our caboose store and noted the nice new storage container for the Diesel dept. To be honest he did seem genuinely sympathetic and did hint that he would talk to management. We called it a day and went home about 3.15.
15 July : posted by Rick
Peter Allan Martin Charles and Rick in today. One of those days where a lot of time and energy exerted but not a lot to show for it. Charles spent the day with Lloyd the electrician checking Norchard in order to prepare an estimate for the compulsory electrical inspection and then they went to Lydney Junction for an estimate for wiring the platform lights.
Peter and Allan were finishing the cables to the new ticket office. Rick and Martin were looking at the SPT at the stop board opposite Norchard signal box. The drop wire had been damaged where it went out of the green conduit and entered the concrete troughing next to the signal box. Working was difficult as trains were being shunted back and forth. A temporary repair was made and the drop wire will have to renewed at a later date. The SPT was checked and proved to have an intermittent handset. This should have been easy to rectify as the whole front plate unplugs and a new one can be plugged in. When we returned with the new front plate we discovered the plugs were of different designs and so would not connect. In the end a new phone was found and the whole thing changed.
Meanwhile Chris Bull was using the DMU to check the line and while at Lydney Junction retrieved our 3 section ladder and delivered it to Norchard and it is now safely in our store.
21 July : posted by Rick
Martin Janet and Rick spent the morning making a temporary repair to the roof of the old DMU container.
22 July : posted by Rick
Rick Martin Allan Peter and Charles started moving stores from the telecoms container to the old DMU container. A good days work was done and we all went home for a rest.
15 July : posted by Rick
Rick Janet and Zara spent the morning moving stores into the old DMU container. There is a lot of electrical gear left in the telecoms container - it needs someone with electrical knowledge to sort it - there is still quite a bit left to move but a couple of mornings hard graft should do it. If anyone is available before next Wednesday could you let me know? All of the DMU stuff has been moved so we have a clear run now. As soon as Chris Bull lets me know when the new container is arriving I will let you know.
26 July : Posted by Rick
Today Peter Rick and Janet moved the last of our stuff into the temporary container and managed to take the wooden shelving down and switch off the electric supply at the fuse box (units covered and labeled). We just need one last push to disconnect the electrical wiring and take the light units down. Anyone available Tuesday from 10ish for an hour or so ? Then the job will be done.
28 July : Posted by Rick
This is an update on what has gone so far. Unfortunately I have been at the DFR most days for the last fortnight and may have the details of times slightly mixed up. Those in today- Allan Charles Martin Peter Rick. Tidying up the old caboosh today. Electric disconnected all associated wiring - lights etc removed. Corrugated metal roof sheets cut away at corners to allow lifting chains access. All old shelving removed and stored near office.
Lydney junction SPT phone reported faulty. Visited PM testing ok!
29 July : Posted by Rick
Those in today - Allan Charles Martin Peter Rick. Man with Lorry and crane on site at 9 o’clock and moving containers all over. Our container disappeared off to Bendalls while we tried to level a new base ready for the new container. Back from Bendalls and lifted our new container onto new base (with a bit of jiggling).
After lunch Peter and Allan went to help John Metherall sort out new wi-fi in his rail coach. Charles Martin and Rick went to Lydney junction to double check the SPTs - all ok except bell not ringing on station tele.
1 August : Posted by Rick
Still no racking but the supplier has said it should be here Monday sometime. Peter Rick Janet Tanya Mick Zara (some of my family) were in today. We fitted the framework ready for the electric supply. Moved some of the stores from the yellow container into the new container and started sorting some of them. Spent most of the rest of the day untangling the blue draw rope that was in the blue bin. It is now all coiled up neatly on drums ready for use.
Sunday is a day of rest - no DFR. Monday I hope to receive the new racking! Tuesday I would be grateful if some of you will be available to assist moving the rest of our stores into the new container.
Then job done, we can get back to normal.
3 August : Posted by Rick
Hurray the racking has arrived - driver drove off. Wait !! There are too many packages and some are for people in Exeter London and Witney. On closer inspection only some of the racking was what I ordered.
Janet Martin and Rick spent the afternoon trying to get some sort of sense out of what we had and managed to get 3 racks built and a lot of bits out of other packs to build a part rack. Went home to recover.
4 August : Posted by Rick
Janet Allan and Rick in today. Allan managed to interpret some of the instructions and with a joint effort we built two more racks. In the middle of this we had Chris Bull checking how far we had got and then the TNT delivery driver arrived again and gave us two more packages but he would not take back the packages we did not want. The new packages did have my name on them so we constructed two more racks. By 2 o’clock we were knackered so Allan went home and Janet and Rick had lunch.
After lunch we had a tidy up and put six packages in the office in the hope that they will be returned. The X team should be in tomorrow ( Rick’s family who are younger and fitter ) so we should get the rest of the stores transferred. I also spoke to Adam Dickinson about putting a duct under the new track before the rails go down. He assured me that it would be done and he was aware of the buried duct on the other side of the track.
5 August : Posted by Rick
Rick Janet Tanya Allan Martin Peter and Charles in today. Peter and Allan cabled and fitted suitable sockets for John Metherall in his sales coach. The rest of us spent the the day completing the move from the yellow container to the new container. Doesn’t sound much but it was very hard work.
During the afternoon David Collins came to see us and asked about the damaged cable at Parkend and asked if it was possible to reconnect it even if only temporary. Rick said that it was possible but we would only go ahead and do it if the DFR honoured what had been said at the board meeting - replacing the cable would be top of the list when funds were available.
We all went home at 3 o’clock.
Later while at home Charles rang Rick to say that Alex Davies had e mailed two pictures of damaged cables at Goatfield to him and would he deal with it. Naturally he forwarded the pictures to Rick. He went out about 6 o’clock to Middle Forge and verified the damage. The two SPT phones and the strowger exten were all working ok so it was decided to have a proper look next week.
6 August : Posted by Rick
Rick Martin Tanya and Mick in today. The original rusty two bay dexion shelving was dismantled and removed from the yellow container. Container now empty !! We all went home about 12 o’clock.
PS - Rick had a phone call from Roger Phelps this afternoon confirming that the DFR will buy new cable as soon as funds allow but in the meantime would we make a temporary joint to enable the token to function.
7 August - 7 September : Posted by Rick
Rick on holiday until 19th Aug - on return started tidying the office and the new container (store). Peter on holiday - Allan injured. Martin and Janet listing all items in the new container - Rick identifying.
Work started to reconnect the token cct between Whitecroft and Parkend and was spaced out over a couple of weeks due to weather and wasp nests. two pairs were proved end to end of the damaged cable ( this section works on bunched pairs) and then a link cable was connected temporarily from each end to the newly fitted cabinets nearby (these were already fitted ready for the new cable length). New pairs testing ok.
The token was tested but did not work. On testing the whole cct - an earth was proved on one leg in the length of cable from DP1 to Parkend exchange (this cable could be going down, we need to check it again later).
Fortunately we were able to divert the token direct to the ground frame. Still not working!! This has taken us at least 36 man hours so far.
9 September : Posted by Rick
Martin and Janet working in caboose listing all stores (this is going to take more than a few hours). Rick working with Roger Phelps on the token. Cct reversed still not working - token machine at Parkend faulty?
David Collins now involved with Roger - fault in token machine cleared - token cct now working.
Framework fitted in container and existing electric cable brought in and fitted to framework.
16 September : Posted by Rick
Martin Janet and Rick in today. All working on listing what is in the caboose.
18 September : Posted by Rick
Peter and Rick in today to meet Steve Robbins - electrician to advise and price job of connecting mains to container.
19 September : Posted by Rick
Received a call from Sam Bolingbroke at 13.00 hrs to report the Goatfield treadle was not working and the cable had been damaged. I visited Goatfield - found the damaged cable and made a temporary repair to give service by 14.45hrs. The cable appears to have been damaged by grass cutting. Chris Bull said none of his men have worked along there for some time. Informed Alastair by e mail and he said he would investigate. Spoke to Steve Harris and he is going to see if he can reroute a new cable.
23 September : Posted by Rick
Peter Martin, Rick and Janet in today. Martin and Janet still recording what is stored in the caboose. Rick is putting up the framework to support the racking and help to fit the lighting to. We were all pleased to see Charles who arrived at midday and we all had lunch together in the Marquee (social distancing). Martin and Rick went with Bernard (museum) to see his electric wiring in the hut behind the Station masters office. He wants our electrician to check and certificate it. Charles has supplied the consumer unit for our container and some cable - Peter has supplied the earth spike and boxes and Rick has supplied the light and cable.
29 September : Posted by Rick
Peter and Rick in today. Steve the electrician arrived - connected everything tested and gave us a certificate. He also checked Bernards.
30 September : Posted by Rick
Martin Janet Rick Allan and Peter in today. Absolutely awful weather today - more like monsoon weather all day. Peter and Allan worked in the office and Rick, Janet and Martin worked in the caboose - good job we had the lights connected yesterday. We all went home early.
7 October : Posted by Rick
Those in today, Rick, Janet, Martin, Peter, Allan, and Charles.
Peter and Allan worked on the digital cabinet in our office fitting jumpers and labels - the raspberry pi is also playing up. Janet and Martin continue collating the items in the new storage container (about 1/3 recorded ).
Charles and Rick had a look at the newly dug trench for the electric supply cable to the coal pile lighting tower and later supplied red conduit and yellow warning tape ready for trench filling. They also liaised and advised the C and W dept and Chris Bull on possible methods of supplying electricity to the new storage containers. Electricity and lighting was connected to the telecoms container last Tuesday and today the safety certificate was given in to the DFR office for their records. A busy day all round but not a lot to show for it.
14 October : Posted by Rick
Rick Janet Peter Martin and Allan in today. Tea coffee and chat to start ( all sitting outside and socially distancing of course ).
Martin andJanet in caboose and recording stores - it seems to be working - two items wanted from the list and quickly found using the numbering system. Rick and Allan managed to put a draw rope through the ducting ready for burial and then provision of telecoms cable to supply comms to the container and the rest of that side of the Norchard site. Peter and Allan then did some more work on the data cabinet.
John and Jean made an unexpected visit to Norchard and wanted to know if we had a spare 50 volt power unit - we had two but we needed one for our future use and the other was not powerful enough for his purposes so unfortunately we were unable to help.
Martin and Janet went home about 1300hrs. The rest had lunch then Peter and Allan took some data parts and an unwanted CC System to Lydney Junction node for storage. Rick had a tidy and locked up and went home.
19 October : Posted by Rick
Rick in attendance today at 11.00 hrs to unlock the office and allow the fire extinguisher to be checked. All ok. Locked up 11.30 hrs. Stay safe.
21 October : Posted by Rick
Rick, Peter, Martin and Janet in today. Weather windy and heavy rain. We took morning tea and coffee in the marquee on the platform to have a team brief to discuss the board meeting on Monday.
At the board meeting last Monday one of the topics on the agenda was the replacement cable between Whitecroft and Parkend. It quickly became obvious that most of the board did not want to spend money on a new cable. Dr Alastair Clarke actually said to the meeting - as the token is working why do we need a new cable. It was then pointed out that we had not officially applied for a PO (purchase order) and that nothing further could be done until we had - but the PO had to have the suppliers name and the price. Unfortunately this is difficult as the price of copper varies so much. It was then decided that we should apply for a PO but not have the price or supplier and the board would then discuss it further.
NOTE : the token was only reconnected after David Collins asked if it was possible. We said only if the board make a commitment to buy the cable when they could afford it and realised that it was a temporary connection as the cable is in poor condition. The next day Roger Phelps sent an e mail confirming that a new cable would be purchased when funds allowed.
Another item that was on the agenda was the fire alarm and who should be responsible for testing it. We used to test it and then some years ago ( at least 4 years ) we passed it to James in the shop. He now claims no memory of this event. As I see it, this is the responsibility of the Admin staff, not a volunteer who may or may not be on site at the time when a test should take place. This has not been a good week for the telecoms dept.
After tea we helped Peter to finish some work in the cabinet in the office. Then because the weather had not improved we all went home at 12.30. Stay well.
25 October : Posted by Rick
Rick Janet and Martin in today. Weather very bad with heavy rain. After tea we sheltered in the caboose and continued recording the stock. Rick checked Norchard exchange - all ok. After lunch we returned to the caboose - very slow going but items are being labelled. Weather still very bad - rain pouring down and thundering so we went home early.
4 November : Posted by Rick
Rick and Peter in today - weather much better today even some sunshine. During tea we discussed the latest flurry of e mails about drainage at Middle Forge/Goatfield. The drainage needs major repairs and will probably require the siding to be lifted and refitted. The existing siding has been laid on top of the S&T and Comms cables which may be a problem. There is talk of moving the cables!
We decided that as it was a nice day we would hitch a ride on the DMU in order to have a look at Middle Forge and refresh our memories of where the DP s were. Chris Bull was the DMU driver and readily agreed to our request but as the train was fully booked he allowed us to ride in the alternative driver’s cab at the rear. This gave us a splendid view of the track and allowed us to see what we wanted. Whilst chatting to Chris Bull we learnt of plans to install a cycle track between Lydney and Parkend ( a Greenway ) and he e mailed me the details. As the plan is to use some railway land we may have more cable problems ahead.
After lunch we decided to have a closer look at why the new clock relay set in the data cabinet in our office was not working. A good bit of old fashioned logical fault finding soon gave us the answer - HR connection inside the cabinet ( krone ). With the offending wire reconnected we had regular clock pulses. We can now tell what time it is in our office. It was decided we should go home while we’re on a high.
11 November : Posted by Rick
Rick and Peter in today. The data cabinet in our office was discussed again.
Then we went to the cafe to decide what to do about the data cabinet in there (fitted when cctv was provided) and is starting to come away from the wall at the top. The main problem of course is that the wall is basically a partition wall with plasterboard facing. Hanging a heavy data cabinet high up is inherently dangerous. The bottom of the cabinet appears to be glued to the plasterboard with 4 plasterboard screws inside the cabinet - at least one of which just goes round and round. We emptied the cabinet to have a better look and try to find a solid fixing. A lot of time was wasted trying to find a vertical baton so that we could make a secure fix. Eventually we decided to put all the cards and patch cables back - although we did reconfigure to our record card. James was asked to test the cafe card machine and it all worked ok. We now have to think of a plan B.
18 November : Posted by Rick
Rick Janet and Peter in today. After tea we all went to the caboose where we had left a small pile of unwanted items under a tarpaulin until we could find a forever home for them. Unfortunately they were now in the way of the work for the new siding. We moved them to alongside our container now that there was space and recovered the old wooden bench too. Janet set to work in the caboose and Peter and Rick returned to the cafe to have another try to find any vertical batons for a fixing for the data cabinet. After lunch we returned to the cafe and used filler to fill any holes in the plasterboard. We have now decided to fit a support shelf beneath the data cabinet. A search for suitable brackets and shelving was successful but would need preparatory work by Peter - so we had a cup of tea and went home.
24 November : Posted by Rick
Rick went to Norchard to meet lineside and P way depts who were going to Middle forge to clear the scrub and lift the siding. Rick was to keep an eye on the cables and check for old disused ones with the help of Bob Hawker. The lineside guys worked like maniacs with their chainsaws and brush cutters but didn’t damage any cables. The P way guys started unbolting the rails and then used the train to drag out the rails. The main telecoms cable was alongside and also underneath the railway sleepers but the P way crew were very careful taking them out and don’t seem to have damaged any of the cable. Five drop wires were recovered during the day and taken back to Norchard.
A return visit will be required to check the layout of cables and DPs ( telecoms and S&T ) when P way have finished lifting all the track and sleepers. A long and tiring day but it was very pleasant working alongside other depts showing that it is possible - we should do it more often!
25 November : Posted by Rick
Rick Janet Peter and Allan in today. Almost like old times with four of us in today. Peter and Allan went to the cafe to see about fitting the new shelf. Janet was back in the caboose. Rick went to meet Steve Harris to discuss the work required at Middle Forge. Allan left shortly before lunch.
After lunch Peter and Rick went to look at the SPT phone at Norchard crossing hut - noisy. As there was no one in the signal box it took us a while to realise it was working but there was a no ring tone! Relay contacts seem ok so could be a faulty relay coil but it seems to test ok. Had a cup of tea and decided to check again next week - it might be a bit warmer.
28 November : Posted by Rick
Rick Janet and Peter in today. We walked the line from Norchard to Goatfield to survey the DPs and cables. We now have a clear idea how the cables run but if the plan is to run all the cables along the line - Middle forge to Town crossing we may have a problem. The telecoms cable from Middle forge to Town crossing was replaced about 2 years ago and was run along the fence line (it used to run trackside). More information required.
2 December : Posted by Rick
Rick Janet and Peter in today. While having tea and chatting, Janet and Peter fitted labels to the krone strips ready to fit later. Peter had brought the shelf unit which will be fitted in the cafe under the sagging cabinet Janet set to painting it. Rick and Peter went round checking the clocks. Some were 10 seconds fast but the rest were all over the place so we had to fetch the ladder from the caboose. We corrected the one in the workshop and gave it a clean - you can now tell the time halfway down the workshop. The clock in the machine shop was corrected and we tried to correct the clock in 9683 but a very thin watchmakers screwdriver is required - next week. The clock in the shop/entrance is a two sided affair and the one facing the incoming customer is ok but the other side is way out. There are bookcases blocking access, so another job for next week.
Some stores were moved from the office to the caboose and the joint boxes that have been laying about were tidied into a waiting space.
We all went to Norchard signal box and the krone strips were fitted in the new box conn ready for the future cable changeover! After a late lunch we all decided that it was very cold and we were no longer young and went home.
End of Year Report : Posted by Rick
It’s been a funny old year and I just want to clarify a few of the events that have happened as I think some of you do not have the full picture.
Sam left at the end of January and although I did try to speak to him - e mail him he did not reply. At a meeting of the Telecoms dept it was agreed that I should take over.
At the end of February I was informed that our steel storage container (caboose) would have to be moved as two new ones were being located next to ours for the DMU group and the Carriage and Wagon group. I spoke to Duncan about this and was told this had been arranged some months previously - the Telecoms dept had shown no interest in a new container ( about £3000 ) and now there was no money or budget for the Telecoms dept.
In March the corona virus arrived and the DFR was closed.
In July preparations were being made to reopen to the public. At this time I met Chris Bull and we discussed cable damage along the railway and I ranted about moving the caboose as it was rusting away and we had missed the chance to get a new one. He seemed genuinely sympathetic and said he would have a word with the Board. A few days later we met again and he said that he could possibly get us a new container if we would pay for the crane to remove the old and deliver the new ( £150 ). This seemed too good to be true but the slush fund was good so I agreed. To cut it short the containers arrived - we moved our stores to a spare and then back into the new one after we had fitted new racking. I still do not know who funded the caboose but it is much appreciated. The slush fund is nearly depleted but I have to thank those who gave generously for the extras - racking and electrics. Martin, Charles, Paul, Rick and Peter. I also need to thank all those whose hard work made it all possible. Rick, Martin, Alan, Peter, Janet, Tanya, Zara, Mick and Chris Bull.
In August, just as we were finishing the move I was approached by David Collins - S and T dept - who asked if it was possible to reconnect the token to Parkend. This had been reconnected temporarily a year ago - after catastrophic damage to the cable between Whitecroft and Parkend but which had broken down again in January and we had decided not to repair it until a new cable was provided. The board had agreed to purchase a new cable when it had funds.
After a discussion I said we would not repair it again as I felt the board would change their mind about a new cable. David disagreed but said he would speak to Roger Phelps (Chairman) about it.
In August Roger Phelps rang me at home and sent me an e-mail confirming that a new cable would be purchased ASAP and asked if I would repair the token in the meantime. Against my better judgment I repaired the token.
At the October Board meeting the purchase of the Whitecroft to Parkend cable was on the agenda but the general feeling of the board was that it was not needed. Dr Alastair Clarke said the token is now working so we do not need a new cable.
In December at a Zoom Board meeting Roger Phelps remonstrated with the board about the Parkend cable saying that there had been an accident involving contractors working at Parkend and there had been difficulties in communication. He then asked the board if they considered it a safe working area. No strong answers. He then suggested we buy a new cable and most appeared to agree. I will be getting quotes when people are back at work and we will have to see what happens.
I now look forward to a better year at the railway. We already have major work to do at Middle Forge and hopefully at Parkend. We have a fistful of minor repairs all over the railway and there has been a total lack of maintenance on the Strowger equipment.
Finally I need to thank all of you for your invaluable help and support in getting us through 2020 and hope that you will continue your support through 2021.
Happy New year to you all and I hope it’s better than 2020.
Once again Happy New Year and above all
Rick G, DFR Telecoms dept
Return to the Diary Menu
Page provided by John Bathgate
This page was last updated on
4th January 2021